OpenStudy (anonymous):

Which expression is equivalent to x6 x2 A x4 x3 B x5 x3 C x7 x3 D x9 x3

OpenStudy (anonymous):

B

OpenStudy (anonymous):

How is it b

OpenStudy (anonymous):

x^6x^2 makes x^8. x^5x^3 also makes x^8

OpenStudy (aum):

\(\Large x^6x^2 = x^{6+2} = x^8\) Do the same with each answer choice and see which one yields \(\Large x^8\).

OpenStudy (anonymous):

yah just add the exponents

OpenStudy (anonymous):

Thank you